If the gas inside the flask is cooled so that its
Answer #1

1 mm = 1 torr
Hence, we will consider presssure in mm Hg units.

797.3 = Patm + (136.4-103.8) \\ 797.3 = Patm + 32.6 \\ Patm = 797.3 - 32.6 \\ Patm = 764.7 mm


The sum of the heights in two arms = 136.4 + 103.8 = 240.2 mm

Now pressure is reduced to 715.7 mm
715.7 = Patm + \text {height difference} \\ 715.7 = 764.7 + \text {height difference} \\ \text {height difference} = 715.7-764.7 = -49
The height of mercury in the open ended arm = \dfrac {240.5}{2}- \dfrac {49}{2} = 95.8 mm


Hence the second option 95.6 mm is correct answer.